1)Two high school classes took the same test. One class of 20 students made an average grade of 80%; the other class of 30 students made an average grade of 70%. What is the average grade for all students in both classes?
2)The marching band of Success University is practicing for a drill. Two hundred members equally spaced form a large circle and each member is numbered consecutively. What is the number of the member who stands directly across from number 9?
3)Mary’s age is twice Tom’s age and one-third of Sally’s age. If the sum of their ages is 99, how old is Mary?
4)If a 5x9x12 rectangular block of wood is painted red and then cut into 540 1x1x1 cubes, how many cubes will have red paint on them?

Answer by checkley71(8403)

(.8*20+.7*30)/50=(16+21)/50 or 37/50=.74 grade for all 50 students.
100+9=109 the number opposite #9 100 is opposite 0 then add 9 to each we get 9 & 109
m+m/2+3m=99 or 4m+m/2=99 or (8m+m)/2=99 or 9m/2=99 or 9m=99*2 or 9m=198 or m=22
thus Mary is 22 Toms 1/2*22=11 & Sally is 3*22=66 thus 22+11+66=99 or 99=99
a 9*12*5 cube has 2 faces 9*12=216 now we have 2 faces 3*9=54 & 2 faces 3*12=72
adding up all these exclusive faces we get 216+54+72=342